12. (5 points) Determine the pressure in a 125-L tank containing 56.2 kg of oxygen gas at 21°C.
Answer #1

given volume - 125L Mall of oxygen go = 562 kg = 56200.999m Malay may z = 15/2 = 32g/mol Temperature = a/c =21+ 273 K = 294 Kpressure = 338.92 atm